# Short-term CO2 emissions forecasting: insight from the Italian electricity market

### v1

@@ 4 Results
Table 1: Summary statistics of carbon emissions by Italian electricity market zone, reported for the training, calibration, and test sets.
@@
$\text{RMSE}=\frac{1}{24}\sum_{h=1}^{24}\text{RMSE}_{h}.$ (12)
$R^{2}=\frac{1}{24}\sum_{h=1}^{24}R^{2}_{\text{h}}.$ (14)
The DM test is used to assess the statistical significance of differences in forecast accuracy between individual models and forecast combinations (Diebold and Mariano,, 1995). This test provides a formal framework to compare the predictive performance of two competing models based on their forecast error differentials. The hypothesis system of the one-sided DM test is defined as follows:
$\begin{cases}H_{0}:\text{Models $i$ and $j$ have the same forecast accuracy;}\\ H_{1}:\text{Model $i$ is less accurate than model $j$ (i.e. it has a significantly higher forecast error).}\end{cases}$
The DM test statistic is calculated as
$DM=\frac{\bar{d}}{\sqrt{(1/T)\sum_{k=-m}^{m}\gamma_{k}}}$ (15)
where: $\bar{d}$ is the sample mean of the loss differential $d_{t,ij}$, defined as $d_{t,ij}=e_{i,t}^{2}-e_{j,t}^{2}$ with $e_{i,t}$ and $e_{j,t}$ denoting the forecast errors of the models $i$ and $j$ at time $t$. The loss function $L(\cdot)$ is the squared error; $\gamma_{k}$ denotes the $k$-th order autocovariance of the loss differential series $d_{t,ij}$; $T$ is the number of forecast periods; $m$ is the maximum lag order used to estimate the long-run variance of $d_{t,ij}$. Under the null hypothesis $H_{0}$, the DM statistic asymptotically follows a standard normal distribution: $DM\sim\mathcal{N}(0,1)$.
4.1 Single model results
In the calibration test, the best performing models are selected by minimising the hourly average RMSE. The selected models are reported in Table 2. For Italy, the selected models are SARIMAX(2,1,0)(1,0,0)24, FARX(2), GAM(3,1), and KNN(10). Table B1 in Appendix B reports the hourly RMSE, hourly average RMSE and R2, while the panel (a) in Figure 6 displays the relative RMSE compared to the Naïve model. The selected GAMs for Sicily and Sardinia are GAM(2,0) and GAM(1,1), respectively. These specifications are more parsimonious compared to GAM(3,1): GAM(2,0) excludes the weekly stochastic component and incorporates autoregressive effects up to 2 days, while GAM(1,1) includes the weekly stochastic component but limits autoregressive effects to 1 day.
For the test set, Table 3 reports the forecast performance of the best individual model of each modelling approach for Italy. RMSE values are provided for each hour of the day (0 to 23), along with the hourly average RMSE and the hourly average out of sample $R^{2}$. The GAM(3,1) model produces the lowest RMSE for most hours, especially during the daytime hours from 6:00 to 23:00. In contrast, the FARX(2) model performs best in the early morning hours from 0:00 to 5:00. The Naïve, TBATS, and SARIMAX(2,1,0)(1,0,0)24 models generally exhibit higher RMSE values. KNN(10) model shows moderate performance but remains consistently behind the FARX(2) and GAM(3,1) models. However, it is interesting to note that during the early morning hours, the KNN(10) model emerges as the second-best performing model. The hourly average RMSE and out-of-sample $R^{2}$ further confirm that the GAM(3,1) model is the best, followed by the FARX(2) and KNN(10) models. Although the GAM(3,1) model provides the most accurate and reliable forecasts during daytime hours, the FARX(2) model may be preferable in the early morning. These findings are further supported by the panel (b) in Figure 6, which displays the relative RMSE compared to the Naïve model for Italy in the test set.
Figures A1-A2 in Appendix A illustrate the hourly performance of the best models for the seven market zones.
Zone SARIMAX SFARX GAM KNN
Italy SARIMAX(2,1,0)(1,0,0)24 FARX(2) GAM(3,1) KNN(10)
North SARIMAX(3,1,0)(1,0,0)24 FAR(2) GAM(3,1) KNN(10)
Centre North SARIMAX(1,1,0)(1,0,0)7 FARX(2) GAM(3,1) KNN(35)
Centre South SARIMAX(3,1,0)(1,0,0)24 FARX(1) GAM(3,1) KNN(10)
South SARIMAX(1,1,0)(1,0,0)24 FARX(3) GAM(3,1) KNN(10)
Calabria SARIMAX(3,1,0)(1,0,0)24 FAR(2) GAM(3,1) KNN(15)
Sicily SARIMAX(2,1,0)(1,0,0)24 FARX(3) GAM(2,0) KNN(5)
Sardinia SARIMAX(1,1,0)(1,0,0)24 FARX(1) GAM(1,1) KNN(10)
@@ 4.1 Single model results
Table 3: Hourly RMSE, hourly average RMSE and R2 in the test set for Italy. The best performance in bold.
Hour Naïve SARIMAX(2,1,0)(1,0,0)24 FARX(2) GAM(3,1) TBATS KNN(10)
0 1747.82 757.95 586.27 2014.15 828.55 895.14
1 1649.77 999.12 724.59 1913.53 1076.61 968.44
2 1576.46 1141.88 801.13 1835.65 1247.85 1010.81
3 1579.05 1224.22 865.96 1787.27 1555.41 1050.16
4 1601.69 1260.28 930.96 1714.19 1969.83 1101.75
5 1801.48 1379.73 1152.19 1572.13 2204.72 1280.53
6 2634.33 2115.97 1927.47 1448.36 2343.35 1899.96
7 3408.24 2891.68 2627.1 1685.3 2999.86 2488.23
8 3643.49 3142.7 2829.13 1777.15 3255.62 2671.11
9 3683.27 3199.33 2892.96 1850.17 3358.22 2722.47
10 3701.08 3224.43 2949.4 1941.42 3457.43 2776.08
11 3650.4 3197.24 2944.01 1985.57 3471.27 2793.27
12 3328.83 2960.48 2727.41 1926.09 3280.71 2616.79
13 3398.8 3023.03 2759.18 1954.43 3350.12 2641.73
14 3650.12 3242.94 2933.18 2017 3641.52 2812.86
15 3745.77 3365.16 3059.69 2005.06 3911.17 2950.08
16 3725.02 3397.52 3039.63 1978.08 4046.44 2975.35
17 3526.17 3302.41 2882.34 1883.04 4003.37 2884.23
18 3311.3 3218.4 2704.5 1791.15 3875.57 2708.54
19 3170.96 3177.96 2615.32 1715.96 3781.29 2614.37
20 3033.02 3070.66 2514.62 1682.37 3533.3 2533.37
21 2683.83 2753.22 2246.8 1619.47 2978.78 2337.93
22 2330.99 2438.37 1979.21 1613.7 2497.61 2150.83
23 2017.95 2245.55 1774.22 1613.48 2166.09 1988.89
Ave. 2858.33 2530.43 2186.14 1805.2 2868.11 2203.04
R2 0.49 0.59 0.68 0.81 0.47 0.69
4.2 Forecast combination results
Table 4: Forecast combinations.
Model Method Combination
COMB1 SARIMAX, SFARX, GAM, TBATS, KNN SA COMB1.SA
BG COMB1.BG
COMB2 SFARX, GAM SA COMB2.SA
BG COMB2.BG
SEL COMB2.SEL
COMB3 SARIMAX, GAM SA COMB3.SA
BG COMB3.BG
COMB4 GAM, KNN SA COMB4.SA
BG COMB4.BG
SEL COMB4.SEL
Each combination combines the best performing models from each class. BG, SA and SEL denote the combination method: BG refers to Bates and Granger, (1969), SA denotes the simple average, and SEL indicates the selection-based combination method. For example, COMB1.BG for Italy combines the forecasts of SARIMAX(2,1,0)(1,0,0)24, FARX(2), GAM(3,1), TBATS and KNN(10) using Bates and Granger, (1969). In contrast, COMB2.SEL combines the forecasts of FARX(2) and GAM(3,1) using the selection-based approach.
The results for Italy are shown in Figure 7 and listed in Table B2 in the Appendix B, which reports the hourly average RMSE of the forecast combinations in the test set. The simple average combination produces a lower RMSE than Bates and Granger, (1969) in most cases, with the exception of combinations involving SFARX and GAM (COMB2.SA) and KNN and GAM (COMB4.SA), where Bates and Granger, (1969) performs better. The best performing combination, COMB2.SEL, achieves an hourly average RMSE of 1564.54, which represents a reduction of 50.4% and 15.1% compared to the Naïve and GAM(3,1) models, respectively. In addition, COMB2.SEL reduces the RMSE of 28.6% compared to the FARX(2) model. According to COMB2.SEL, for the early hours of the day ($0$ 05:00), the forecasts of FARX(2) are used, while for the remaining hours (06:00 - 23:00), the predictions of GAM(3,1) are applied. COMB2.SEL reduces the RMSE of 28.6% compared to the FARX(2) model. As an alternative, COMB4.SEL, the second-best forecast combination, achieves an hourly average RMSE of 1616.44, representing a reduction of 48.7% compared to the Naïve model. COMB2.BG, the third-best forecast combination, reduced the hourly average RMSE of 47% compared to the Naïve model.
Figure 7: Relative RMSE of forecast combinations for Italy, compared to the Naïve model.
Table 5 presents the p-values of the DM test for pairwise comparisons of forecast accuracy among individual models and forecast combinations for Italy. Each value represents the p-value of the DM test that compares the model in the row ($i$) with the model in the column ($j$). In terms of single models, the SARIMAX(2,1,0)(1,0,0)24 model is significantly less accurate than all other single models and forecast combinations. Similarly, the TBATS model is significantly less accurate than all other models and forecast combinations, except the SARIMAX(2,1,0)(1,0,0)24 model. The FARX(2) model outperforms the SARIMAX(2,1,0)(1,0,0)24 and TBATS models. The KNN(10) model performs better than the SARIMAX(2,1,0)(1,0,0)24, FARX(2), and TBATS models. However, the forecast accuracy of the FARX(2) and KNN(10) models is almost the same, as indicated by the p-value of 0.0394. This suggests that at the 1% significance level, the FARX(2) and KNN(10) models have equivalent forecast performance. The GAM(3,1) model outperforms all individual models, but is less accurate than the forecast combinations. This aligns with the RMSE results, confirming that forecast combinations enhance accuracy beyond the best single models, such as GAM(3,1).
SARIMAX(2,1,0)(1,0,0)24 FARX(2) GAM(3,1) TBATS KNN10 COMB2.BG COMB2.SEL COMB4.SEL
SARIMAX(2,1,0)(1,0,0)24 - $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$
FARX(2) 1.$0$ $<0.05$ 1.0000 0.0394 $<0.05$ $<0.05$ $<0.05$
GAM(3,1) 1.0000 1.$0$ 1.0000 1.0000 $<0.05$ $<0.05$ $<0.05$
TBATS 1.0000 $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$
KNN10 1.0000 0.9606 $<0.05$ 1.$0$ $<0.05$ $<0.05$ $<0.05$
COMB2.BG 1.0000 1.0000 1.0000 1.0000 1.$0$ $<0.05$ $<0.05$
COMB2.SEL 1.0000 1.0000 1.0000 1.0000 1.0000 1.$0$ 1.0000
COMB4.SEL 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 $<0.05$
For the seven market zones, Figures A3-A4 display the relative RMSE compared to the Naïve model, while the hourly average RMSE is reported in Table B3 and the p-values of the DM test are reported in Tables B4-B10 in the Appendix B.
The results of the DM test are consistent with those obtained using the RMSE metric, reinforcing the conclusion that forecast combinations enhance the predictive accuracy of short-term carbon emissions compared to single models. Among the combination methods, COMB2.SEL emerges as the best performer for Italy as a whole, as well as for the North, Centre-North, Centre-South, South and Calabria. In most of these cases, COMB4.SEL or COMB2.BG are the second-best combinations. Among individual forecasting models, GAM(3,1) consistently delivers the best overall performance. However, for Sicily, COMB2.BG provides the most accurate forecast combination, with FARX(3) identified as the best single model. In the case of Sardinia, COMB1.BG is the best performing combination, while FARX(1) is the best single model.
5 Conclusions
This study investigates the short-term forecasting of carbon emissions from electricity generation in the Italian power market. Using hourly data from 2021 to 2023, several statistical models and forecast combination methods are evaluated and compared at the national and zonal levels. Four main model classes are considered: linear parametric models, functional parametric models, non-parametric and possibly non-linear models, and a semi-functional approach. Forecast combinations include the simple averaging, Bates and Granger, (1969), and the selection-based approach.
For mainland Italy, including the national level, North, Centre-North, Centre-South, South, and Calabria, GAM provides the most accurate forecasts during daytime hours, whereas functional parametric models perform best during the early morning period. In general, GAM emerges as the most effective individual model based on the hourly average RMSE and DM test results. Among the forecast combination methods, the selection-based approach consistently outperforms all individual models and combination methods, achieving a significant reduction in RMSE and frequently emerging as the top performer in the DM test.
In Sicily, the functional parametric model outperforms all other individual models across all hours. Furthermore, the combination of GAM forecast and the functional parametric model with Bates and Granger, (1969) consistently produces the best results, outperforming both individual models and alternative combination strategies.
In Sardinia, all individual models show reasonably good forecast performance (Figure A4). However, the functional parametric model once again stands out as the best performing individual model and also achieves the best performance among all model combinations in terms of the hourly average RMSE. According to the DM test, combinations that include all individual models with the approach of Bates and Granger, (1969) produce superior performance.
These results underscore the importance of selecting an appropriate method for the prediction of short-term CO2 emissions. Once again, the findings highlight how forecast performance can vary significantly between different areas of the electricity market, reflecting differences in local energy mixes. In this study, for example, the results obtained for Italy and its mainland regions differ from those observed in the two islands Sicily and Sardinia. These differences can be attributed to the different energy mixes and the level of interconnection with the mainland, both of which influence electricity generation and related electricity generation (Sapio and Spagnolo,, 2016, 2020; Weinhold and Mieth,, 2021). In addition, the use of forecast combinations, which integrate the strengths of multiple individual models, has been shown to improve predictive accuracy in most cases, particularly when the models capture different structural features of the data (Wang et al.,, 2023).
For policy makers, these findings emphasise the need to take into account regional specificities and to rely on robust ensemble-based forecasting tools when designing and implementing emission reduction strategies (Abolghasemi et al.,, 2025).
This work can be expanded in several ways. First, future work can integrate other approaches and models, such as neural networks and deep learning methods (Lowry,, 2018; Bokde et al.,, 2021; Maji et al., 2022b, ). Second, multi-day forecasts can also be explored, which can potentially allow for more advanced demand-side management strategies (Maji et al., 2022a, ; Mason et al.,, 2018; Girolimetto et al.,, 2024). Third, it is worth investigating the forecast reconciliation methods (Wickramasuriya et al.,, 2019; Di Fonzo and Girolimetto,, 2024; Girolimetto and Di Fonzo,, 2024), as they provide a coherent framework to ensure consistency between forecasts at different levels of aggregation. This is particularly important in the context of electricity-related CO2 emissions, where forecasts may be required at the zonal and national levels.
Acknowledgments
Data availability statement
The authors do not have permission to share the data.
Appendix A Figures
Figure A3: Relative RMSE of forecast combinations for North, Centre-North, Centre-South, and South, compared to the Naïve model. Test set.
Figure A4: Relative RMSE of forecast combinations for Calabria, Sicily and Sardinia, compared to the Naïve model. Test set.
Appendix B Tables
Table B1: Hourly RMSE, hourly average RMSE and R2 in the calibration set for Italy. The best performance in bold.
Hour Naïve SARIMAX(2,1,0)(1,0,0)24 FARX(2) GAM(3,1) TBATS KNN(10)
0 1934.4 906.31 726.32 1901.91 1165.37 1047.4
1 1850.62 1204.73 852.96 1839.2 1547.37 1131.61
2 1787.36 1356.43 893.17 1744.02 1803.55 1130.84
3 1732.41 1457.05 939.54 1657.98 2201.58 1161.25
4 1730.57 1485.27 979.79 1609.97 2599.61 1170.56
5 1993.67 1628.97 1219.38 1488.41 2738.53 1384.42
6 3019.32 2434.23 2032.11 1495.9 2792.9 2139.71
7 3802.94 3218.2 2623.42 1816.09 3302.3 2664.07
8 4006.97 3413.01 2768.9 1944.74 3461.96 2759.48
9 4054.56 3441.76 2810.67 1984.83 3510.49 2775.11
10 4032.99 3410.96 2834.07 2011.23 3544.32 2792.73
11 3984.19 3400.23 2840.02 2022.3 3554.66 2779.49
12 3859.39 3335.71 2810.18 2006.73 3526.03 2745.07
13 3912.7 3404.01 2857.08 2047.56 3666.78 2832.36
14 4081.13 3553.39 2954.62 2106.45 3991.52 2966.27
15 4137.96 3650.6 2992.64 2143.02 4191.56 3000.32
16 4059.9 3646.91 2975.58 2100.58 4289.79 2967.97
17 3761.57 3505.27 2777.73 1946.52 4145.6 2804.65
18 3451.93 3345.27 2584.59 1809.17 3918.66 2628.55
19 3282.8 3277.99 2487.29 1722.14 3762.69 2539.81
20 3212.19 3223.67 2416.48 1703.04 3590.9 2525.25
21 2987.37 3001.79 2278.99 1701.77 3272.09 2443.42
22 2642.74 2684.89 2055.81 1731.99 2854.17 2265.98
23 2315.08 2493.59 1901.82 1669.91 2541.37 2114.49
Ave. 3151.45 2770.01 2192.22 1841.89 3165.58 2282.12
R2 0.36 0.49 0.67 0.79 0.36 0.66
Table B2: Hourly RMSE, hourly average RMSE and R2 in the test set for Italy. The best performance in bold.
Hour COMB1.BG COMB2.SA COMB2.BG COMB4.SA COMB5.BG COMB6.SA COMB2.SEL COMB4.SEL COMB9.BG COMB10.SA
0 692.5 675.93 1289.24 1097.32 711.81 1075.58 586.27 895.14 1298.32 1217.07
1 899.86 753.97 1228.41 1056.15 914.03 1059.51 724.59 968.44 1244.59 1173.12
2 1021.43 792.23 1181.9 1024.44 1033.69 1041.42 801.13 1010.81 1195.64 1130.9
3 1162.58 846.46 1165.87 1022 1105 1036.39 865.96 1050.16 1170.88 $1$.51
4 1318.9 914.11 1121.55 992.32 1129.82 993.82 930.96 1101.75 1138.49 1087.43
5 1450.16 1044.79 1050.86 964.88 1233.44 961.33 1152.19 1280.53 1110.77 1082.93
6 1947.89 1598.52 1263.71 1310.53 1944.2 1388.02 1448.36 1448.36 1363.55 1391.2
7 2662.76 2241.52 1737.8 1842.86 2701.19 2010.84 1685.3 1685.3 1811.36 1861.02
8 2895.44 2441.46 1893.5 2011.24 2945.46 2213.26 1777.15 1777.15 1955.59 2010.65
9 2961.97 2505.28 1964.83 2080.83 3004.51 2283.41 1850.17 1850.17 2010.88 2064.47
10 3010.29 2563.1 2039.8 2150.89 3035.17 2342.19 1941.42 1941.42 2072.37 2123.47
11 2997.83 2563.02 2063.17 2167.88 3012.28 2343.89 1985.57 1985.57 2096.17 2145.5
12 2781.64 2368.46 1934.23 2019.86 2784.41 2169.44 1926.09 1926.09 1970.96 2012.98
13 2828.5 2405.45 1968.13 2054.03 2846.05 2223.26 1954.43 1954.43 2004.71 2046.4
14 3046.2 2578.15 2080.13 2179.45 3057.98 2387.77 2017 2017 2130.14 2178.56
15 3212.84 2697.81 2126.29 2242.63 3171.34 2454.38 2005.06 2005.06 2191.02 2248.93
16 3287.64 2718.88 2107.52 2224.87 3197.25 2452.34 1978.08 1978.08 2199.11 2260.25
17 3229.02 2612.82 1997.41 2107.76 3096.89 2335.23 1883.04 1883.04 2111.04 2172.6
18 3137.57 2474.64 1884.32 1984.75 3007.14 2225.4 1791.15 1791.15 1984.86 2041.31
19 3075.76 2383.58 1803.4 1902.62 2960.8 2154.27 1715.96 1715.96 1887.62 1943.27
20 2921.66 2273.66 1744.21 1835.42 2857.85 2074.03 1682.37 1682.37 1824.59 1877.43
21 2525.83 1985.1 1583.96 1649.94 2556.65 1857.56 1619.47 1619.47 1678.8 1723.54
22 2156.21 1737.75 1470.61 1502.92 2263.54 1679.01 1613.7 1613.7 1564.53 1597.46
23 1906.19 1575.25 1397.58 1403.28 2087.3 1585.06 1613.48 1613.48 1481.82 1503.73
Ave. 2380.44 1948 1670.77 1701.2 2360.74 1847.81 1564.54 1616.44 1729.08 1750.24
R2 0.63 0.75 0.83 0.82 0.64 0.79 0.85 0.84 0.82 0.81
Table B3: Hourly average RMSE on the test set across market zones and forecast combinations. The best performance in bold.
Model North Centre-North Centre-South South Calabria Sicily Sardinia
COMB1.SA 1224.54 107.10 451.54 315.85 323.72 198.96 115.04
COMB1.BG 1494.22 114.06 491.77 340.87 362.85 208.69 113.93
COMB2.SA 1033.12 103.97 428.53 307.12 300.83 200.12 121.98
COMB2.BG 1008.39 104.31 438.47 309.97 302.92 197.41 117.41
COMB2.SEL 930.16 104.23 417.97 303.55 296.76 207.65 126.44
COMB3.SA 1144.31 107.57 429.08 305.63 314.58 202.06 118.90
COMB3.BG 1453.90 118.94 497.15 349.70 371.58 221.66 116.91
COMB4.SA 1082.17 104.31 441.82 317.17 309.00 208.21 128.90
COMB4.BG 1064.92 104.81 448.99 322.93 312.69 209.10 126.66
COMB4.SEL 962.92 105.53 426.29 311.95 307.95 212.98 130.41
Table B4: Diebold-Mariano test p-values for pairwise comparisons for North.
SARIMAX(3,1,0)(1,0,0)24 FAR(2) GAM(3,1) TBATS KNN(10) COMB2.BG COMB2.SEL COMB4.SEL
SARIMAX(3,1,0)(1,0,0)24 - $<0.05$ $<0.05$ 1.0000 $<0.05$ $<0.05$ $<0.05$ $<0.05$
FAR(2) 1.$0$ $<0.05$ 1.0000 1.0000 $<0.05$ $<0.05$ $<0.05$
GAM(3,1) 1.0000 1.$0$ 1.0000 1.0000 $<0.05$ $<0.05$ $<0.05$
TBATS $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$
KNN(10) 1.0000 $<0.05$ $<0.05$ 1.$0$ $<0.05$ $<0.05$ $<0.05$
COMB2.BG 1.0000 1.0000 1.0000 1.0000 1.$0$ $<0.05$ $<0.05$
COMB2.SEL 1.0000 1.0000 1.0000 1.0000 1.0000 1.$0$ 1.0000
COMB4.SEL 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 $<0.05$
@@ Appendix B Tables
SARIMA(1,1,0)(1,0,0)7 FARX(2) GAM(3,1) TBATS KNN(35) COMB2.BG COMB2.SEL COMB4.SEL
SARIMA(1,1,0)(1,0,0)7 - $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$
FARX(2) 1.$0$ 0.1485 1.0000 0.9997 $<0.05$ $<0.05$ $<0.05$
GAM(3,1) 1.0000 0.8515 - 1.0000 0.9980 $<0.05$ $<0.05$ $<0.05$
TBATS 1.0000 $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$ $<0.05$
KNN(35) 1.0000 $<0.05$ $<0.05$ 1.$0$ $<0.05$ $<0.05$ $<0.05$
COMB2.BG 1.0000 1.0000 1.0000 1.0000 1.$0$ 0.4611 0.9583
COMB2.SEL 1.0000 1.0000 1.0000 1.0000 1.0000 0.5389 - 1.0000
COMB4.SEL 1.0000 1.0000 1.0000 1.0000 1.0000 $<0.05$ $<0.05$

### v2

@@ Appendix B Tables
Table B1: Summary statistics of carbon emissions by Italian electricity market zone, reported for the training, calibration, and test sets.
@@
Table B2: Hourly RMSE, hourly average RMSE and R2 in the test set for Italy. The best hourly average performance in bold.
Hour Naïve SARIMAX(2,1,0)(1,1,1)24 FARX(2) GAM(3,1) TBATS KNN(10)
0 1747.82 603.62 586.27 2014.15 828.55 895.14
1 1649.77 772.52 724.59 1913.53 1076.61 968.44
2 1576.46 887.19 801.13 1835.65 1247.85 1010.81
3 1579.05 979.06 865.96 1787.27 1555.41 1050.16
4 1601.69 1039.11 930.96 1714.19 1969.83 1101.75
5 1801.48 1197.70 1152.19 1572.13 2204.72 1280.53
6 2634.33 1985.94 1927.47 1448.36 2343.35 1899.96
7 3408.24 2739.02 2627.1 1685.3 2999.86 2488.23
8 3643.49 2964.60 2829.13 1777.15 3255.62 2671.11
9 3683.27 3026.55 2892.96 1850.17 3358.22 2722.47
10 3701.08 3079.08 2949.4 1941.42 3457.43 2776.08
11 3650.4 3065.49 2944.01 1985.57 3471.27 2793.27
12 3328.83 2830.48 2727.41 1926.09 3280.71 2616.79
13 3398.8 2888.15 2759.18 1954.43 3350.12 2641.73
14 3650.12 3075.39 2933.18 2017 3641.52 2812.86
15 3745.77 3174.88 3059.69 2005.06 3911.17 2950.08
16 3725.02 3158.19 3039.63 1978.08 4046.44 2975.35
17 3526.17 2995.49 2882.34 1883.04 4003.37 2884.23
18 3311.3 2803.59 2704.5 1791.15 3875.57 2708.54
19 3170.96 2717.13 2615.32 1715.96 3781.29 2614.37
20 3033.02 2628.65 2514.62 1682.37 3533.3 2533.37
21 2683.83 2376.60 2246.8 1619.47 2978.78 2337.93
22 2330.99 2136.75 1979.21 1613.7 2497.61 2150.83
23 2017.95 1956.45 1774.22 1613.48 2166.09 1988.89
Ave. 2858.33 2295.07 2186.14 1805.2 2868.11 2203.04
R2 0.49 0.65 0.68 0.81 0.47 0.69
Table B3: Hourly RMSE, hourly average RMSE and R2 in the calibration set for Italy. The best hourly average performance in bold.
Hour Naïve SARIMAX(2,1,0)(1,1,1)24 FARX(2) GAM(3,1) TBATS KNN(10)
0 1934.4 783.63 726.32 1901.91 1165.37 1047.4
1 1850.62 999.65 852.96 1839.2 1547.37 1131.61
2 1787.36 1082.1 893.17 1744.02 1803.55 1130.84
3 1732.41 1185.39 939.54 1657.98 2201.58 1161.25
4 1730.57 1234.81 979.79 1609.97 2599.61 1170.56
5 1993.67 1435.93 1219.38 1488.41 2738.53 1384.42
6 3019.32 2299.66 2032.11 1495.9 2792.9 2139.71
7 3802.94 2994.56 2623.42 1816.09 3302.3 2664.07
8 4006.97 3148.54 2768.9 1944.74 3461.96 2759.48
9 4054.56 3183.03 2810.67 1984.83 3510.49 2775.11
10 4032.99 3177.17 2834.07 2011.23 3544.32 2792.73
11 3984.19 3154.29 2840.02 2022.3 3554.66 2779.49
12 3859.39 3094.95 2810.18 2006.73 3526.03 2745.07
13 3912.7 3164.33 2857.08 2047.56 3666.78 2832.36
14 4081.13 3297.02 2954.62 2106.45 3991.52 2966.27
15 4137.96 3334.79 2992.64 2143.02 4191.56 3000.32
16 4059.9 3317.88 2975.58 2100.58 4289.79 2967.97
17 3761.57 3145.27 2777.73 1946.52 4145.6 2804.65
18 3451.93 2966.86 2584.59 1809.17 3918.66 2628.55
19 3282.8 2921.25 2487.29 1722.14 3762.69 2539.81
20 3212.19 2887.36 2416.48 1703.04 3590.9 2525.25
21 2987.37 2772.98 2278.99 1701.77 3272.09 2443.42
22 2642.74 2531.22 2055.81 1731.99 2854.17 2265.98
23 2315.08 2333.09 1901.82 1669.91 2541.37 2114.49
Ave. 3151.45 2518.57 2192.22 1841.89 3165.58 2282.12
R2 0.36 0.58 0.67 0.79 0.36 0.66
Table B4: Hourly RMSE, hourly average RMSE and R2 in the test set for Italy. The best hourly average performance in bold.
Hour Naïve COMB1.SA COMB1.BG COMB2.SA COMB2.BG COMB2.SEL COMB3.SA COMB3.BG COMB3.SEL COMB4.SA COMB4.BG COMB4.SEL
0 1747.82 661.87 604.23 1097.32 1289.24 586.27 1060.16 582.88 603.62 1217.07 1298.32 895.14
1 1649.77 722.76 766.5 1056.15 1228.41 724.59 1012.8 721.35 772.52 1173.12 1244.59 968.44
2 1576.46 751.29 869.63 1024.44 1181.9 801.13 970.41 815.59 887.19 1130.9 1195.64 1010.81
3 1579.05 799.52 992.32 1022 1165.87 865.96 955.49 896.18 979.06 1111.51 1170.88 1050.16
4 1601.69 864.62 1127.82 992.32 1121.55 930.96 921.9 947.2 1039.11 1087.43 1138.49 1101.75
5 1801.48 1011.65 1303.43 964.88 1050.86 1152.19 907.31 1094.69 1197.7 1082.93 1110.77 1280.53
6 2634.33 1601.77 1941.65 1310.53 1263.71 1448.36 1351.69 1863.71 1448.36 1391.2 1363.55 1448.36
7 3408.24 2237.78 2654.42 1842.86 1737.8 1685.3 1941.4 2600.42 1685.3 1861.02 1811.36 1685.3
8 3643.49 2432.35 2873.78 2011.24 1893.5 1777.15 2119.54 2820.3 1777.15 2010.65 1955.59 1777.15
9 3683.27 2496.06 2939.24 2080.83 1964.83 1850.17 2184.94 2882.21 1850.17 2064.47 2010.88 1850.17
10 3701.08 2555.58 2997.02 2150.89 2039.8 1941.42 2244.29 2934.31 1941.42 2123.47 2072.37 1941.42
11 3650.4 2555.23 2988.63 2167.88 2063.17 1985.57 2243.89 2921.13 1985.57 2145.5 2096.17 1985.57
12 3328.83 2354.56 2763.59 2019.86 1934.23 1926.09 2073.19 2693.34 1926.09 2012.98 1970.96 1926.09
13 3398.8 2392.11 2818.71 2054.03 1968.13 1954.43 2120.25 2749.74 1954.43 2046.4 2004.71 1954.43
14 3650.12 2558.27 3019.46 2179.45 2080.13 2017 2261.93 2931.77 2017 2178.56 2130.14 2017
15 3745.77 2670.54 3152.49 2242.63 2126.29 2005.06 2316.39 3025.95 2005.06 2248.93 2191.02 2005.06
16 3725.02 2680.88 3174.75 2224.87 2107.52 1978.08 2308.07 3011.56 1978.08 2260.25 2199.11 1978.08
17 3526.17 2566.7 3051.27 2107.76 1997.41 1883.04 2187.39 2855.74 1883.04 2172.6 2111.04 1883.04
18 3311.3 2416.13 2885.09 1984.75 1884.32 1791.15 2047.19 2671.57 1791.15 2041.31 1984.86 1791.15
19 3170.96 2326.35 2800.29 1902.62 1803.4 1715.96 1959.16 2584.74 1715.96 1943.27 1887.62 1715.96
20 3033.02 2231.57 2682.26 1835.42 1744.21 1682.37 1890.89 2498.68 1682.37 1877.43 1824.59 1682.37
21 2683.83 1971.42 2374.18 1649.94 1583.96 1619.47 1706.85 2254.15 1619.47 1723.54 1678.8 1619.47
22 2330.99 1745.92 2090.49 1502.92 1470.61 1613.7 1555.33 2023.42 1613.7 1597.46 1564.53 1613.7
23 2017.95 1587.52 1876.65 1403.28 1397.58 1613.48 1463.18 1853.92 1613.48 1503.73 1481.82 1613.48
Ave. 2858.33 1924.69 2281.16 1701.2 1670.77 1564.54 1741.82 2176.44 1581.96 1750.24 1729.08 1616.44
R2 0.49 0.76 0.66 0.82 0.83 0.85 0.81 0.69 0.85 0.81 0.82 0.84
Table B5: Diebold-Mariano test p-values for pairwise comparisons in the North. H0 equal forecast accuracy; H1: the column model outperforms the row model.
SARIMAXa FAR(2) GAM(3,1) TBATS KNN(10) COMB2.SEL COMB3.SEL
SARIMAXa - <0.05<0.05 <0.05<0.05 1.00 <0.05<0.05 <0.05<0.05 <0.05<0.05
FAR(2) 1.00 - <0.05<0.05 1.00 1.00 <0.05<0.05 <0.05<0.05
GAM(3,1) 1.00 1.00 - 1.00 1.00 <0.05<0.05 <0.05<0.05
TBATS <0.05<0.05 <0.05<0.05 <0.05<0.05 - <0.05<0.05 <0.05<0.05 <0.05<0.05
KNN(10) 1.00 <0.05<0.05 <0.05<0.05 1.00 - <0.05<0.05 <0.05<0.05
COMB2.SEL 1.00 1.00 1.00 1.00 1.00 - 1.00
COMB3.SEL 1.00 1.00 1.00 1.00 1.00 <0.05<0.05 -
  • Note. a the model order is (3,1,0)(1,1,1)24.
@@
SARIMAa FARX(2) GAM(3,1) TBATS KNN(35) COMB2.SA COMB2.SEL
SARIMAa - <0.05<0.05 <0.05<0.05 <0.05<0.05 <0.05<0.05 <0.05<0.05 <0.05<0.05
FARX(2) 1.00 - 0.15 1.00 1.00 <0.05<0.05 <0.05<0.05
GAM(3,1) 1.00 0.85 - 1.00 1.00 <0.05<0.05 <0.05<0.05
TBATS 1.00 <0.05<0.05 <0.05<0.05 - <0.05<0.05 <0.05<0.05 <0.05<0.05
KNN(35) 1.00 <0.05<0.05 <0.05<0.05 1.00 - <0.05<0.05 <0.05<0.05
COMB2.SA 1.00 1.00 1.00 1.00 1.00 - 0.88
COMB2.SEL 1.00 1.00 1.00 1.00 1.00 0.12 -
  • Note. a the model order is (1,1,0)(1,1,0)24.